In the basic Keynesian model of income determination, calculate the expenditure multiplier, given the following consumption function:

C = 200 + 0.50Y


1
Expert's answer
2021-05-24T12:37:43-0400

Given that,

C = 200 + 0.50Y

Expenditure Multiplier = "\frac{1}{1-MPC}" ,

MPC = 0.50

"\therefore" Multiplier = "\frac{1}{1-0.50}"

="\frac{1}{0.50}"

= 2